OpenStudy (anonymous):

Write the decimal as a percent 0.9

OpenStudy (anonymous):

Just times it by 100, and get 90. That's the answer. 90%

OpenStudy (akashdeepdeb):

We know 5% = 5/100 and 20% = 20/100 and so on. So x% = x/100 = 0.9 Thus x = 100*0.9 = 90 Thus 0.9 = 90% = 90/100 = 0.9. Understood? :)

OpenStudy (goformit100):

A warm Welcome to OpenStudy. I can help guide you through this useful site. You can ask your questions to me or you can message me. Please use the chat for off topic questions. Remember to give the person who helped you a medal by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it.